OpenStudy (anonymous):

What quadrant is 11pi/3 in? Find sin, cos, and tan theta.

OpenStudy (anonymous):

is there a way to do this besides looking at the unit circle because I am confused on where to find 11pi/3

OpenStudy (anonymous):

@Jaynator495 @taylor12344 @ParthKohli

jaynator495 (jaynator495):

im sorry i dont know how to explain it without giving the answer away, and as thats not allowed ill try and call someone who can help more with this :)

jaynator495 (jaynator495):

@just_one_last_goodbye @TheSmartOne

OpenStudy (aum):

\[ \frac{11\pi}{3} = \frac{9\pi+2\pi}{3} = 3\pi + \frac 23\pi = 2\pi + \pi + \frac 23 \pi \]

jaynator495 (jaynator495):

figures, the one i didnt call but htought of came and helped lol

OpenStudy (aum):

\[2\pi + \pi + \frac 23\pi\]\(2\pi\) is one full circle and can be ignored. \[\pi + \frac 23 \pi = \pi + \frac 12 \pi + \frac 16\pi\] In which quadrant does that lie?

OpenStudy (anonymous):

@aum do u know

OpenStudy (anonymous):

dont struggle with it, its fine guys, i gtg anyway. aum u can always comment here or message me if u wanna help me out, i appreciate it

OpenStudy (aum):

\[ \pi + \frac 12 \pi + \frac 16\pi \\ \pi = 180 \text{ degrees} \\ 180 + 90 + 30 = 300 \text{ degrees}. \]

OpenStudy (aum):

Which quadrant is 300 degrees?

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

i mean 4 srry

OpenStudy (aum):

Yes.

OpenStudy (anonymous):

Do i find sin cos and tan theta by looking at the parenthesis on the unit circle

OpenStudy (aum):

If radians is confusing, convert it to degrees and it will be easier to figure out the quadrant.

OpenStudy (anonymous):

i need to figure out the next part now ^^

OpenStudy (aum):

The first value within the parenthesis, that is the x-coordinate is the cosine value and the y-coordinate is the sine value.

OpenStudy (anonymous):

can u check my answers for that

OpenStudy (anonymous):

cos= -1/2 sin= (-sqrt 3)/2 tan= sqrt 3 pretty sure the last one is right, and maybe the sin

OpenStudy (aum):

cosine is positive in the fourth quadrant and so it should be 1/2 (not -1/2). tangent is negative in the fourth quadrant.

OpenStudy (aum):

\[ \cos\left( \frac{11\pi}{3} \right) = \frac 12 \\ \sin\left( \frac{11\pi}{3} \right) = -\frac {\sqrt{3}}{2} \\ \tan\left( \frac{11\pi}{3} \right) = -\sqrt{3} \\ \]

OpenStudy (anonymous):

thanks!!! @aum

OpenStudy (aum):

yw.